# Self-Guided Open-Vocabulary Semantic Segmentation

### 07 Dec 2023 v1

@@ Abstract
Vision-Language Models (VLMs) have emerged as promising tools for open-ended image understanding tasks, including open vocabulary segmentation. Yet, direct application of such VLMs to segmentation is non-trivial, since VLMs are trained with image-text pairs and naturally lack pixel-level granularity. Recent works have made advancements in bridging this gap, often by leveraging the shared image-text space in which the image and a provided text prompt are represented. In this paper, we challenge the capabilities of VLMs further and tackle open-vocabulary segmentation without the need for any textual input. To this end, we propose a novel Self-Guided Semantic Segmentation (Self-Seg) framework. Self-Seg is capable of automatically detecting relevant class names from clustered BLIP embeddings and using these for accurate semantic segmentation. In addition, we propose an LLM-based Open-Vocabulary Evaluator (LOVE) to effectively assess predicted open-vocabulary class names. We achieve state-of-the-art results on Pascal VOC, ADE20K and CityScapes for open-vocabulary segmentation without given class names, as well as competitive performance with methods where class names are given. All code and data will be released.
@@ 1 Introduction
Semantic segmentation is a computer vision task that involves grouping pixels into coherent regions based on their semantic categories. While notable advancements have been made, existing semantic segmentation models primarily rely on labeled datasets with predefined categories, limiting their ability to handle unknown classes. In contrast, humans possess an open-vocabulary understanding of scenes, recognizing thousands of distinct categories. Recent studies have focused on the development of segmentation models to meet similar capabilities [21, 18], with models that leverage Vision-Language Models (VLMs) as an emerging category [14, 10, 7, 16, 5, 36, 3, 34, 35]. Such VLMs learn rich multi-modal features from billions of image-text pairs.
Figure 1: Self-Guided Open-Vocabulary Semantic Segmentation. Our method is capable of determining and segmenting relevant open vocabulary class names automatically.
Given the scale, training VLMs comes at a extremely high computational cost. Because of this, it is common in literature to build upon pre-trained VLMs such as CLIP [22] or BLIP [15]. However, applying these VLMs on per-pixel tasks is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image. Previous attempts have employed a two-stage approach where masked regions are fed to the VLMs to produce CLIP embeddings, which are subsequently correlated with text encodings of text prompts to perform classification [33, 16, 7]. Other works directly align per-pixel embeddings with text embeddings retrieved from CLIP [14] based on text prompts provided by the user at test time. While training these models, the text prompts usually correspond to the ground truth class names of the dataset in question. In this paper, we take the capabilities of VLMs one step further and tackle a more challenging task in which the relevant class names are automatically generated. We enable self-guided open-vocabulary semantic segmentation, in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user or predefined class names.
To tackle this task, we introduce a Self-Guided Semantic Segmentation (Self-SEG) framework. Our framework can adapt to any pre-trained segmentation backbone limited by the need for textual input describing the class names to be potentially segmented. In order to automate this, we propose BLIP-Cluster-Caption (BCC) which consists of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network without any additional supervised training.
Experimental evaluations on Pascal VOC [9], ADE20K [37] and CityScapes [6] showcase the effectiveness of our framework. Self-SEG achieves state-of-the-art results on self-guided open-vocabulary semantic segmentation and competitive results with methods where class names are given. It is capable of predicting crisp segmentation masks without having any information a priori about which objects to segment (see Fig. 1). For effective evaluation on public datasets, we furthermore propose an LLM-based Open-Vocabulary Evaluator (LOVE), tasked with mapping open-vocabulary predictions to class names in the dataset using a Large Language Model (LLM). All code will be publicly released to ensure reproducibility and to facilitate further research.
In summary, our contributions are as follows: 1) we introduce Self-SEG, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from clustered BLIP embeddings; and 3) we propose LOVE, a novel evaluation method for open-vocabulary semantic segmentation which utilizes an LLM.
Through qualitative and quantitative analyses, we demonstrate the effectiveness of our Self-Guided Semantic Segmentation framework on multiple public datasets.
@@
Figure 2: Different Semantic Segmentation Tasks in Comparison. Left: In traditional semantic segmentation, an image is segmented into fixed, predefined dataset classes used during annotation. Middle: In open-vocabulary segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented. Right: In self-guided segmentation, relevant object categories are generated from the image automatically, enabling a true open-vocabulary understanding of scenes without a human in the loop.
@@ 2 Related Work
Open-Vocabulary Segmentation (OVS). OVS refers to the challenging task of segmenting images based on a list of arbitrary text queries. Contrary to traditional segmentation methods, OVS is not limited to a predefined set of classes. The field of OVS has seen significant advancements through different approaches. Approaching OVS as a zero-shot task, pioneering work by ZS3Net [1] and SPNet [29] trained custom modules that translate between visual and linguistic embedding spaces and segment by comparing word vectors and local image features. These methods attempt to learn the challenging connection between two very different and independent latent spaces. Consequently, the field of OVS has seen the introduction of methods incorporating Vision-Language Models (VLMs). VLMs pre-train a visual feature encoder with a textual feature encoder on large amounts of image-text pairs, enabling image and text representations to reside in the same feature space. One of such VLMs, Contrastive Language-Image Pre-training (CLIP) [22], has triggered a wave of segmentation approaches that harness pretrained CLIP embeddings for OVS. LSeg [14] proposed a Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments. More popularly, two-stage approaches such as OpenSeg [10], OPSNet [3], OVSeg [16], ZSSeg[33] and POMP [24] first obtain class-agnostic masks and use CLIP to classify each mask afterwards. Other works, including MaskCLIP [7], FC-CLIP [36], and SAN [34], focus on utilizing intermediate representations from a frozen CLIP encoder. For example, MaskCLIP uses the token features of the ViT-based CLIP encoder and compares local features to class label embeddings instead of pooling them into a global CLIP embedding. The remarkable diversity in the usage of CLIP for OVS extends to works which leverage relationships between image semantics and class labels in CLIP space [5], use CLIP for binary mask prediction [35] or ensure segmentation consistency at different granularity levels using CLIP [28]. Several works have proposed to tackle OVS outside the realm of VLMs using text-to-image diffusion models [11, 32] or custom transformer networks [30, 17, 31, 38]. Most notably, X-Decoder introduces a generalized transformer model capable of performing various vision-language tasks using a single set of parameters. One common and important limitation of the described methods is the need for textual input as a form of guidance by the user. Our work presents a novel framework which enables self-guidance by automatically finding relevant object categories using localized image captioning. Closest and concurrent to our work is the Zero-Guidance Segmentation paradigm [25], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user. Despite the similar setting, there are key differences. While we use the BLIP [15] model for both image clustering and caption generation, they design a complex pipeline consisting of image clustering with DINO, obtaining vision-language embeddings with CLIP and a custom attention-masking module, to then caption these with CLIP-guided GPT-2. This requires converting between three different latent representations throughout the process. In contrast, we achieve superior segmentation quality by enabling self-guidance using any OVS model.
@@
Image Captioning with Image-Text Embeddings. Image Captioning is the task of describing the content of an image using natural language. A number of approaches have been proposed for the task, including ones utilizing VLMs. Approaches such as CLIP-Cap [20] and CLIP-S [4] utilize CLIP embeddings to guide text generation. Other approaches rely on BLIP [15], a VLM in which the task of decoding text from image features was one of the tasks of the pretraining process. This enables BLIP to perform image captioning without additional components. Furthermore, the text decoder is guided by local patch embeddings in a way that enables local captioning based on a specific area. This unique capability, discovered during the development of our method, emerges organically in BLIP despite being trained solely with image-level captions.
@@
Visual Phrase Grounding. Visual phrase grounding aims to connect different entities mentioned in a caption to corresponding image regions [26]. This task resembles OVS, as it aims to find correspondences between text and image regions. However, predicted areas do not have to precisely outline object boundaries and can be overlapping. Visual grounding has been approached in a self-guided manner, in which heatmaps of regions of the image are generated from image-level CLIP embeddings, which are then captioned by the BLIP encoder [27]. One disadvantage of this approach is that the region determined by CLIP can indicate multiple objects at once, hence multiple objects are captioned for that region, preventing pixel-level class-specific predictions. In our work, BLIP is used both as an encoder and decoder, which enables for category-specific regions and captioning.
@@
Figure 3: Method Overview. BLIP encodings are clustered, aligned and denoised before being decoded into nouns. Generated nouns serve as self-guidance to an OVS model, which outputs the final mask. During evaluation, we process the output mask with our custom evaluator LOVE, which maps predicted nouns to dataset class names.
@@ 3 Method
The aim of this paper is to perform open vocabulary semantic segmentation without training, finetuning or textual input. To this end, we propose to identify relevant object categories in the image by captioning local regions in the image at different scales (Sect. 3.1), filtering the generated captions into meaningful semantic descriptions (Sect. 3.2) and using these as self-guidance for a segmentor (Sect. 3.3). An overview of our method is depicted in Fig. 3.
@@ 3.1 Local Region Captioning
To identify relevant object categories in the image, we employ Bootstrapped Language Image Pretraining (BLIP) [15]. BLIP is a powerful VLM able to perform accurate and detailed image captioning. The BLIP image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text representations reside. Subsequently, the set of embeddings is passed through transformer layers to capture contextual information across the image. Finally, all contextualized embeddings are passed to the decoder to generate text descriptions. Though effective in describing salient objects in the image, captioning models generally fail to describe all elements of a scene in a comprehensive way. This is likely due to the structure of training data, in which image descriptions mostly focus on salient foreground elements. This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-vocabulary segmentation, in which many objects in the image go unnoticed. As a solution to this limitation, we seek to enable BLIP’s text decoder to consider all encoded areas in the image. Therefore, we propose to cluster the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively. Finally, each BLIP feature cluster is fed to the text decoder separately. This enables generating a description of each semantically meaningful area in the image, resulting in a more complete description of the image overall.
Clustering. By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$. We additionally process the image at a larger resolution $\mathbf{X}_{L}\in\mathbb{R}^{512\times 512\times 3}$, as common datasets often have images with higher resolution. The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{L}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\mathbf{B}$ at the two resolutions:
@@
$\mathbf{P}^{\mathbf{X}_{L}}$ =$=\left\{\mathbf{X}^{L}_{ij}\,|\,\mathbf{X}^{L}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 32\right\}$ (2)
$\mathbf{B}^{\mathbf{X}_{D}}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}^{\mathbf{X}_{D}})\big{)}\quad\textbf{B}^{\mathbf{X}_{L}}=T\big{(}f_{\text{MLP}}(\mathbf{P}^{\mathbf{X}_{L}})\big{)}$ (3)
$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{i}$ $=\mathbf{B}^{\mathbf{X}_{R}}_{i}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{i},\quad\forall i\in\{1,2,\ldots,N\}$ (4)
where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{L}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches for the default and large resolutions $D$ and $L$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP:
$T_{l}(X)$ $=\textsc{LayerNorm}\big{(}X+\textsc{MHA}(X,X,X)\big{)}$ (5)
$T^{\prime}_{l}(X)$ $=\textsc{LayerNorm}\big{(}T_{l}(X)+f_{\text{MLP}}(T_{l}(X))\big{)}$ (6)
$T(X)$ $=T^{\prime}_{L}\circ T^{\prime}_{L-1}\circ\ldots\circ T^{\prime}_{1}(X).$ (7)
In Eq. (4), we additionally concatenate a sin-cos positional embedding [^1] Sin-cos positional embeddings are obtained using the positional-encodings Python library. $\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$. Next, we cluster the patches in $\hat{\mathbf{B}}$ for each resolution $R$ using $k$-means clustering [19]:
@@
where $C_{k}$ is the set of $k$ clusters and $\mu_{i}$ are the cluster centroids. Running the clustering procedure with two unique values for $k$ on two different image resolutions results in four different cluster assignments.
Cross-clustering Consistency.
Each run of $k$-means clustering labels its clusters independently from others, yielding a correspondence issue between clusters across runs. To resolve this, we relabel the cluster indices to a common reference frame with the following steps:
  1. 1.
    We pick the set of clusters with the most clusters after $k$-means as a reference set $S$. This is not necessarily the set with highest initial k, because some clusters can become empty during the $k$-means iterations. The reference set determines the indices used for all other sets of clusters $\mathcal{C}$, each with its number of clusters denoted by $|C_{i}|$:
@@
  2. 2.
    Sets of clusters are aligned to the reference set using Hungarian matching [13]. Given the reference set $S$ and another set of clusters $C$, we calculate pairwise IoU (Intersection over Union) between the clusters from $R$ and $C$. Then, each cluster from $C$ is assigned a new index, matching the cluster with the highest IoU from $S$:
@@
    Assign $\text{Assign }c_{j}\text{ to }r_{i}\text{ where }i=\!\underset{i\in\{1,\ldots,|S|\}}{\mathrm{argmax}}\,\text{IoU}(c_{j},r_{i})$ (10)
  3. 3.
    With the labeled sets of clusters, we assign a probability distribution over the clusters to each image patch. For a given patch $p$, let $L(p)=\{L_{1}(p),L_{2}(p),\ldots,L_{m}(p)\}$ be the set of labels assigned to $p$ by the $m$ different sets of clusters. The probability $P(n|p)$ of $p$ being assigned to a particular cluster $n$ is defined as the relative frequency of $n$ among labels $L(p)$:
    $P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }L_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (11)
@@
Cluster Denoising. To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [^2] We use the crfseg Python library. [12] and majority filter. CRF is a discriminative statistical method that is used to denoise predictions based on local interactions between them. In our case, the predictions are a 2D grid of cluster assignment probabilities of the image patches. The CRF implementation we use is specifically tailored for refining 2D segmentation maps. It uses a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices. Key to this process is the use of a Gaussian filter in the pairwise potentials, which ensures spatial smoothness and consistency in the segmentation.
As shown in Fig. 3, the application of the CRF yields clusters which are cleaner and more cohesive than the original aligned $k$-means result. To address any remaining noise in the resulting clusters, we apply a neighborhood majority filter as a final step. For each image patch, we consider the set of patches $\mathcal{N}(i,j)$ in its square neighborhood: $\mathcal{N}(i,j)=\{(i+\delta_{i},j+\delta_{j})\,|\,\delta_{i},\delta_{j}\in\{-1,0,1\}\}$. We calculate the mode value from the cluster indices in that neighborhood and assign it as the new index of the central patch:

### 20 Mar 2024 v2

# Auto-Vocabulary Semantic Segmentation

@@ Abstract
Open-ended image understanding tasks gained significant attention from the research community, particularly with the emergence of Vision-Language Models. Open-Vocabulary Segmentation (OVS) methods are capable of performing semantic segmentation without relying on a fixed vocabulary, and in some cases, they operate without the need for training or fine-tuning. However, OVS methods typically require users to specify the vocabulary based on the task or dataset at hand. In this paper, we introduce Auto-Vocabulary Semantic Segmentation (AVS), advancing open-ended image understanding by eliminating the necessity to predefine object categories for segmentation. Our approach, AutoSeg, presents a framework that autonomously identifies relevant class names using enhanced BLIP embeddings, which are utilized for segmentation afterwards. Given that open-ended object category predictions cannot be directly compared with a fixed ground truth, we develop a Large Language Model-based Auto-Vocabulary Evaluator (LAVE) to efficiently evaluate the automatically generated class names and their corresponding segments. Our method sets new benchmarks on datasets such as PASCAL VOC and Context, ADE20K, and Cityscapes for AVS and showcases competitive performance to OVS methods that require specified class names. All code is publicly available here.
Keywords:
Auto-Vocabulary Semantic Segmentation Open-Vocabulary Semantic Segmentation Semantic Segmentation
@@ 1 Introduction
Semantic segmentation is a computer vision task that involves grouping pixels into coherent regions based on their semantic categories. While notable advancements have been made, existing semantic segmentation models primarily rely on labeled datasets with predefined categories, limiting their ability to handle unknown classes. In contrast, humans possess an open-ended understanding of scenes, recognizing thousands of distinct categories. Recent studies have focused on the development of segmentation models to meet similar capabilities [24, 20], with models that leverage Vision-Language Models (VLMs) as an emerging category [15, 10, 7, 17, 5, 40, 3, 37, 39]. Such VLMs learn rich multi-modal features from large numbers of image-text pairs.
@@
Figure 1: Different Semantic Segmentation Tasks in Comparison. In traditional Semantic Segmentation, an image is segmented into fixed, predefined dataset classes used during annotation. In Open-Vocabulary Segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented. In Auto-Vocabulary Segmentation, relevant object categories are generated from the image automatically, enabling a true open-ended understanding of scenes without a human in the loop.
Figure 2: AutoSeg Exemplary Results. While established segmentation datasets have a fixed set of annotation categories, our method is able to identify and segment object categories beyond the fixed-set ground truth, such as cub, dachshund or pagoda. Images are from the PASCAL VOC [9] and ADE20K [41] datasets.
Given the scale, training VLMs comes at a extremely high computational cost. Because of this, it is common to build upon pre-trained VLMs such as CLIP [25] or BLIP [16]. However, applying these VLMs on per-pixel tasks is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image. Previous attempts have employed a two-stage approach where masked regions are fed to the VLMs to produce CLIP embeddings, which are subsequently correlated with text encodings of text prompts to perform classification [38, 17, 7]. Other works directly align per-pixel embeddings with text embeddings retrieved from CLIP [15] based on text prompts provided by the user at test time. While training these models, the text prompts usually correspond to the ground truth class names of the dataset in question. In this paper, we take the capabilities of VLMs one step further and tackle a more challenging task in which the relevant class names are automatically generated. We enable Auto-Vocabulary Semantic Segmentation (AVS), in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user, predefined class names, additional data, training or fine-tuning.
To tackle this task, we introduce the AutoSeg framework. Our framework can adapt to any Open-Vocabulary Segmentation (OVS) backbone. We propose BLIP-Cluster-Caption (BCC) to generate relevant class names, consisting of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network.
Experimental evaluations on PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] showcase the effectiveness of AutoSeg, achieving state-of-the-art results on AVS and results competitive to OVS methods which are - unlike AutoSeg- limited by the need for pre-specified class names. It is capable of predicting crisp segmentation masks without having any information a priori about which objects to segment (see Fig. 2). For effective evaluation on public datasets, we furthermore propose a Large Language Model-based Auto-Vocabulary Evaluator (LAVE), tasked with mapping auto-vocabulary predictions to fixed ground truth class names in the dataset. All code will be publicly released to ensure reproducibility and to facilitate further research into AVS.
In summary, our contributions are as follows: 1) we introduce AutoSeg, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from enhanced BLIP embeddings; and 3) we propose LAVE, a novel evaluation method for auto-vocabulary semantic segmentation which utilizes an LLM. Through qualitative and quantitative analyses, we demonstrate the effectiveness of our framework for Auto-Vocabulary Semantic Segmentation on multiple public datasets.
@@ 2 Related Work
Open-Vocabulary Segmentation (OVS). OVS refers to the challenging task of segmenting images based on a list of arbitrary text queries. Contrary to traditional segmentation methods, OVS is not limited to a predefined set of classes. The field of OVS has seen significant advancements through different approaches. Approaching OVS as a zero-shot task, pioneering work by ZS3Net [1] and SPNet [33] trained custom modules that translate between visual and linguistic embedding spaces and segment by comparing word vectors and local image features. These methods attempt to learn the challenging connection between two very different and independent latent spaces. Consequently, the field of OVS has seen the introduction of methods incorporating Vision-Language Models (VLMs). VLMs pre-train a visual feature encoder with a textual feature encoder on large amounts of image-text pairs, enabling image and text representations to reside in the same feature space. One of such VLMs, Contrastive Language-Image Pre-training (CLIP) [25], has triggered a wave of segmentation approaches that harness pretrained CLIP embeddings for OVS. LSeg [15] employed the Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments. More popularly, two-stage approaches such as OpenSeg [10], OPSNet [3], OVSeg [17], ZSSeg[38] and POMP [27] first obtain class-agnostic masks and use CLIP to classify each mask afterwards. Other works, including MaskCLIP [7], FC-CLIP [40], and SAN [37], focus on utilizing intermediate representations from a frozen CLIP encoder. For example, MaskCLIP uses the token features of the ViT-based CLIP encoder and compares local features to class label embeddings instead of pooling them into a global CLIP embedding. The remarkable diversity in the usage of CLIP for OVS extends to works which leverage relationships between image semantics and class labels in CLIP space [5], use CLIP for binary mask prediction [39] or ensure segmentation consistency at different granularity levels using CLIP [31]. Several works have proposed to tackle OVS outside the realm of VLMs using text-to-image diffusion models [11, 34] or custom transformer networks [35, 19, 36, 42]. Most notably, X-Decoder introduces a generalized transformer model capable of performing various vision-language tasks using a single set of parameters. One common and important limitation of the described methods is the need for textual input as a form of guidance by the user. Our work presents a novel framework which enables self-guidance by automatically finding relevant object categories using localized image captioning. Closest and concurrent to our work is the Zero-Guidance Segmentation paradigm [28], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user. Despite the similar setting, there are key differences. While we use the BLIP [16] model for both image clustering and caption generation, they design a complex pipeline consisting of image clustering with DINO, obtaining vision-language embeddings with CLIP and a custom attention-masking module, to then caption these with CLIP-guided GPT-2. This requires converting between three different latent representations throughout the process. In contrast, we achieve superior segmentation quality by enabling self-guidance using any out-of-the-box OVS model.
@@
Image Captioning with Image-Text Embeddings. Image Captioning is the task of describing the content of an image using natural language. A number of approaches have been proposed for the task, including ones utilizing VLMs. Approaches such as CLIP-Cap [22] and CLIP-S [4] utilize CLIP embeddings to guide text generation. More recently, the Large Language and Vision Assistant (LLaVA) was introduced with similar capabilities by connecting the visual encoder of CLIP with a language decoder, as well as training on multimodal instruction-following data [18]. Other approaches rely on BLIP [16], a VLM in which the task of decoding text from image features was one of the tasks of the pretraining process. This enables BLIP to perform image captioning without additional components. Furthermore, the text decoder is guided by local patch embeddings in a way that enables local captioning based on a specific area. This unique capability, discovered during the development of our method, emerges organically in BLIP despite being trained solely with image-level captions.
@@
Visual Phrase Grounding. Visual phrase grounding aims to connect different entities mentioned in a caption to corresponding image regions [29]. This task resembles OVS, as it aims to find correspondences between text and image regions. However, predicted areas do not have to precisely outline object boundaries and can be overlapping. Visual grounding has been approached in a self-guided manner, in which heatmaps of regions of the image are generated from image-level CLIP embeddings, which are then captioned by the BLIP encoder [30]. One disadvantage of this approach is that the region determined by CLIP can indicate multiple objects at once, hence multiple objects are captioned for that region, preventing pixel-level class-specific predictions. In our work, BLIP serves both as encoder and decoder, enabling category-specific regions and captioning.
@@
Figure 3: Method Overview. BLIP encodings are clustered, aligned and denoised before being decoded into nouns. Generated nouns serve as self-guidance to a segmentor, which predicts the final mask. When evaluating (purple), our custom evaluator LAVE processes the output, mapping predicted nouns to the fixed-vocabulary annotations.
@@ 3 Method
This work aims to perform auto-vocabulary semantic segmentation without any additional data, training, finetuning or textual input. To this end, we propose to identify relevant object categories in the image by captioning local image regions at different scales (Sect. 3.1), filtering the generated captions into meaningful semantic descriptions (Sect. 3.2) and using the generated classes names as target categories for a segmentor (Sect. 3.3). An overview is depicted in Fig. 3.
@@ 3.1 Local Region Captioning
To identify relevant object categories in the image, we employ Bootstrapped Language Image Pretraining (BLIP) [16]. BLIP is a powerful VLM capable of performing accurate and detailed image captioning. The BLIP image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text representations reside. Subsequently, the set of embeddings is passed through transformer layers to capture contextual information across the image. Finally, all contextualized embeddings are passed to the decoder to generate text descriptions. Though effective in describing salient objects in the image, captioning models like BLIP generally fail to describe all elements of a scene comprehensively. This is likely due to the descriptions in the training data, which mostly focus on salient foreground objects. This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-ended recognition, in which many objects in the image go unnoticed. As a solution to this limitation, we propose BLIP-Cluster-Caption (BCC) to enable the BLIP text decoder to consider all encoded areas in the image. BCC clusters the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively. Finally, each BLIP feature cluster is fed to the BLIP text decoder separately. This enables generating a description of each semantically meaningful area in the image, resulting in a more comprehensive and accurate description of the image overall.
Clustering. By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$. Since common datasets often have images with higher resolution, we additionally process the image at a larger resolution $\mathbf{X}_{H}\in\mathbb{R}^{512\times 512\times 3}$. The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{H}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\hat{\textbf{B}}=\{\hat{\textbf{B}}^{\mathbf{X}_{D}},\hat{\textbf{B}}^{\mathbf{X}_{H}}\}$ at the two resolutions:
@@
$\mathbf{P}^{\mathbf{X}_{H}}$ =$=\left\{\mathbf{X}^{H}_{ij}\,|\,\mathbf{X}^{H}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 32\right\}$ (2)
$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{n}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}_{n}^{\mathbf{X}_{R}})\big{)}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{n},\quad\forall n\in\{1,2,\ldots,N\}$ (3)
where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{H}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches in resolutions $D$ and $H$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP, sequentially propagated with the function composition operator $\circ$:
$T_{l}(X)$ $=X+\textsc{MHA}(\textsc{LayerNorm}(X,X,X)$ (4)
$\hat{T}_{l}(X)$ $=T_{l}(X)+f_{\text{MLP}}(\textsc{LayerNorm}(T_{l}(X)))$ (5)
$T(X)$ $=(\hat{T}_{L-1}\circ\hat{T}_{L-2}\circ\ldots\circ\hat{T}_{0})(X).$ (6)
In Eq. (3), we additionally concatenate a sinusodial positional encoding [^1] Obtained with the positional-encodings Python library. [32] $\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$. Next, we cluster the patches in $\hat{\mathbf{B}}$ for each resolution $R$ using $k$-means clustering [21]:
@@
where $C_{k}$ is the set of $k$ clusters and $\mu_{i}$ are the cluster centroids. Running the clustering procedure with a 2 to 8 range for $k$ on two different image resolutions results in 14 different cluster assignments.
Cross-clustering Consistency. Each run of $k$-means clustering labels its clusters independently from others, yielding a correspondence mismatch between clusters across runs. To resolve this, we relabel the cluster indices to a common reference frame with the following steps:
1. Select $C$ with the most clusters after $k$-means as a reference set $S$. As some clusters end up empty during the $k$-means iterations, this is not always the set with highest initial $k$. The reference set determines the indices used for all other sets of clusters $\mathcal{C}$, each with its number of clusters denoted by $|C_{i}|$:
@@
2. Sets of clusters are aligned to the reference set using Hungarian matching [14]. We calculate pairwise Intersection over Union (IoU) between the clusters from $S$ and $C$. Then, each cluster from $C$ is assigned a new index, matching the cluster with the highest IoU from $S$:
@@
Assign index $\text{Assign index }i\text{ to }c_{j}\text{ where }i=\!\underset{i\in\{1,\ldots,|S|\}}{\mathrm{argmax}}\,\text{IoU}(c_{j},s_{i})$ (9)
3. With the labeled sets of clusters, a probability distribution over the clusters is assigned to each image patch. For a given patch $p$, let $\mathbf{L}(p)=\{\mathbf{L}_{1}(p),\mathbf{L}_{2}(p),\ldots,\\ \mathbf{L}_{m}(p)\}$ be the set of labels assigned to $p$ by the $m$ different sets of clusters. The probability $P(n|p)$ of $p$ being assigned to a particular cluster $n$ is defined as the relative frequency of $n$ among labels $\mathbf{L}(p)$:
$P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }\mathbf{L}_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (10)
@@
Cluster Denoising. To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [^2] We use the crfseg Python library. [13] and majority filter. CRF is a discriminative statistical method that is used to denoise predictions based on local interactions between them. In our case, the predictions are a 2D grid of cluster assignment probabilities of the image patches. Our implementation is specifically tailored for refining 2D segmentation map, using a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices. Key to this process is the use of a Gaussian filter in the pairwise potentials, which ensures spatial smoothness and consistency in the segmentation.
The application of the CRF yields embeddings which are less noisy and more cohesive than the original aligned $k$-means result. To address remaining noise in the embeddings, a neighborhood majority filter is applied as a final step. For each image patch, we consider the set of patches $\mathcal{N}(i,j)$ in its square neighborhood: $\mathcal{N}(i,j)=\{(i+\delta_{i},j+\delta_{j})\,|\,\delta_{i},\delta_{j}\in\{-1,0,1\}\}$. The mode value from the cluster indices in that neighborhood is calculated and assigned as the new index of the central patch: